40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Message ID: 1205     Entry time: Fri Feb 20 05:22:19 2015
Author: Zach 
Type: DailyProgress 
Category: Laser 
Subject: Error calibration -> actuation TFs and new laser frequency noise measurement 

Using the REFL PDH setup I built the other day (and that was detailed somewhat by Nic and Chris W. in CRYO:1204), I calibrated the error response so that I could make some further measurements. To refresh, this is using 30-MHz sidebands applied using one of our fiber phase modulators, sensing with a 1611 in reflection. The sideband drive was 0 dBm.

Using the sidebands as a reference, I calculated the slope at 6.4 nV/Hz:

Note that the error signal is slightly asymmetric, but there is no large offset.

With this information, I made some measurements:

Actuation transfer functions

I wanted to measure the actuation transfer functions afforded by:

  1. The standalone ThorLabs laser diode driver (LDC201C) that is currently used to drive the west laser
  2. The diode driver in the integrated ThorLabs laser controller (ITC502) that is currently used to drive the east laser
  3. Simply driving into the bias tee on the diode

All measurements were done with the west laser head, driving as described and reading out at the error point, then correcting for the loop gain and calibrating to Hz.

Everything is more or less as expected:

  • The LDC201C only claims 3 kHz bandwidth, which is actually a bit of a stretch as usual
  • The ITC502 claims 500 kHz, and this is also a stretch (there is -45° at 40 kHz), but it's not so bad for some early locking
  • Directly driving into the diode works across this measurement band and likely far, far beyond
  • There are some common features, prominently at low frequencies and somewhat so at higher ones that likely arise from the current -> frequency response of the diode itself

 

Laser frequency noise

I now have another calibrated measure of laser frequency noise, wherever it dominates over the PMC length noise. I measured the error signal, corrected for the loop gain and calibrated to Hz. For comparison, I've added the measurement using the Zurich PLL on the beat between the two free-running lasers on 12/17/2014 (see CRYO:1185), as well as the RIO spec for this laser.

As you can see, tonight's measurement agrees quite well with the earlier one upt to ~1 kHz, above which the old measurement is probably marred by the relatively low-bandwidth PLL. It seems that the PMC is quiet enough to see the laser noise throughout, and the new measurement now sits closer to the spec up to the highest available point at 10 kHz. Below ~50 Hz, we are probably seeing the well-documented excess noise from the ThorLabs driver. Everything looks as expected.

Locking via laser feedback

Relatively early in the night, after having measured the actuation transfer functions, I sucessfully locked the cavity via feedback to the laser (as opposed to the PMC PZT) for the first time. Below is a comparison of the OLTFs for a 1-kHz loop using the same servo shape (a pole at 1 Hz) using both actuation schemes.

Because of what I had hooked up at the time, I only did this with the (low-bandwidth) LDC201C, so, while the absence of a ~10-kHz resonance is clear, the phase margin is not improved at all (worsened, actually). I only report this as a milestone, and the margin afforded by the ITC502 or by directly driving via the bias tee should be far better.

ELOG V3.1.3-